OpenStudy (anonymous):

A square poster has sides measuring 2 feet less than the sides of a square sign. If the difference between their areas is 12 square feet. Find the lengths of the sides of the poster and the sign

OpenStudy (anonymous):

It's easy! what is the formula for a square?

OpenStudy (anonymous):

I won't be able to help you if you will just be silent!

OpenStudy (anonymous):

I got 16ft for the sign and 4 ft for the poster does that seem correct to you?

OpenStudy (anonymous):

does it match yours>?

OpenStudy (anonymous):

Well, let's see! a ( is a poster ) and b ( sign) Area = a^2 Area = b^2 b^2 - a^2 = 12 if a = b- 2 we just plug it in our formula! b^2 - (b-2)^2 = 12 b^2 - b^2 + 4b - 4 = 12 4b - 4 = 12 4b = 16 b = 4 Area for b = 16 Area for a = 4 You are correct :)

OpenStudy (anonymous):

yay thanks for checking with me hugs